OpenStudy (anonymous):

Eduardo wants to make 17.1 fl. oz. of a 60% sugar solution by mixing together a 80% sugar solution and a 50% sugar solution. How much of each solution must he use?

hero (hero):

x + y = 17 .50x + .80y = .60(17)

hero (hero):

Can you solve it from here or no?

OpenStudy (anonymous):

no

hero (hero):

x + y = 17 .50x + .80y = 10.2 y = 17 - x y = (10.2 - .50x)/.80 y = y .80(17 - x) = 10.2 - .50x 13.6 - .80x = 10.2 - .50x 13.6 - 10.2 = -.50x + .80x 3.4 = .30x 3.4/.30 = x 11 and 1/3 = x 5 and 2/3 = y

OpenStudy (anonymous):

it's alright. I now see what I have to do. thank you.

OpenStudy (anonymous):

Your x and y solutions look OK.

hero (hero):

nice. I guess the point 1 didn't matter that much